# Towards Relaxed Multimodal Inputs for Gait-based Parkinson’s Disease Assessment

### v1

@@ V-B Parkinson’s Disease Multimodal Datasets
We conducted evaluations on three publicly available multimodal PD gait datasets, summarized statistically in Table IV. All datasets contain gait data from two to three complementary modalities and accompanied by clinically annotated ratings. These characteristics make them suitable for the evaluation of our proposed framework.
FOG Dataset [72] contains video and IMU recordings from 35 PD subjects performing turning-in-place at 30 Hz (video) and 128 Hz (IMU) with clinical FoG annotations. During preprocessing, we first removed irrelevant video segments, then applied MMPose [76] (MotionBERT trained on Human3.6M) to lift 2D detections to anatomically consistent 3D poses. Given the camera framing, we focus analysis on seven lower-limb joints. Each sequence is hip-centered, normalized, and uniformly sampled to F=101 frames. IMU trials are parsed into 6 channels, downsampled 5 times, uniformly split into 36 segments, synchronized to corresponding skeleton segments, and cleaned by discarding inadequate or misaligned segments.
FBG Dataset [73] comprises synchronized full-body kinematics and kinetics from 26 PD subjects walking overground in ON/OFF medication states, including 3D motion capture at 150 Hz and force-plate data, with MDS-UPDRS III and Hoehn–Yahr labels. Skeletons start as 44 keypoints per frame and are interpolated to 17 standard joint centers, normalized, downsampled to 30 Hz, and fixed to (F=101) frames with 51-D per frame. Due to missing or corrupted raw trial-level sensors, we use the dataset’s preprocessed, time-normalized ensemble vGRF signals that average gait cycles per side. During training, we sample either the left or right vGRF per participant as the sensor input of length (101,3).
WearGait Dataset [74] provides synchronized IMU, sensorized insole, and walkway data from 98 people with PD and 83 age-matched controls, collected across multiple sites with clinical metadata. IMUs capture acceleration, angular velocity, magnetic field, and orientation; insoles provide 16-sensor plantar pressures plus embedded inertial signals; all streams are aligned to a pressure walkway reference. For preprocessing, we first resampled each modality from its native 100 Hz export to 30 Hz, preserving temporal alignment. Missing values introduced by device dropouts and alignment padding were filled by mean imputation on a per-channel basis. We then segmented each continuous recording into fixed-length, non-overlapping windows to form training samples.
@@ V-C Overall Evaluation
We present the performance of our proposed framework (TRIP) under three complementary evaluation scenarios to thoroughly validate its effectiveness. The single-modality oracles (e.g., Skeleton Only) establish an upper bound for each modality, and the metric quantifies how multimodal methods perform relative to their strongest unimodal counterpart (negative indicates improvement, positive indicates performance drop).
Asynchronous Training Results. We first conduct experiments under asynchronous and naturally-imbalanced input condition, which happens more frequently in real-clinic settings, to demonstrate that our framework can effectively train with asynchronous modalities, whereas existing multimodal solutions fail due to their reliance on synchronous inputs. Furthermore, we show that our framework even has the potential to outperform oracle models in certain scenarios.
As shown in Table III, across all three datasets, TRIP is consistently best and typically by a large margin over fusion baselines of matched capacity. On the FOG, TRIP attains the top average accuracy of 56.68% which is about 16 pp above the best conventional fusion. Per-modality gains are substantial: Skeleton rises from 47.73% (unimodal) to 63.61% (+15.9 pp) and IMU from 33.47% to 49.76% (+16.3 pp), yielding a large negative =40.97. On the FBG, TRIP again leads with 59.48%. Skeleton improves from 41.8% to 51.24% (+9.4 pp) and vGRF from 60.13% to 67.71% (+7.6 pp), with =17.59. Finally, on the WearGait, despite three modalities, TRIP reaches 71.39% average—best overall—while each single-stream head is competitive (Walkway 71.08%, Insole 63.03%, IMU 80.07%). The aggregate improvement across datasets over unimodal or simple fusions confirms that despite training with relaxed, time-misaligned inputs, our model can still learn complementary subject-invariant information.
Synchronous Training Results. We also conduct experiments using synchronous training inputs to demonstrate that even under the synchronous input setting, our framework maintains superior performance (see Table V). On the FOG, it attains 62.35%, exceeding the strongest unimodal oracle (57.49%) by a large margin and yielding a negative =9.46. On the WearGait, TRIP reaches 84.18% with =28.17, outperforming all fusion baselines (e.g., Early Fusion 81.88%, Cross Attention 79.30%). We omit synchronous evaluation on the FBG because trial-level sensor recordings are largely missing or corrupted and only subject-level, time-normalized vGRF is available, which prevents fair per-trial temporal alignment and underscores the fragility of sync-dependent methods.
Class-Balanced Training for baselines. To ensure that performance gains are not solely due to class imbalance, we repeat the evaluation with class-balanced training, using a naïve inverse-frequency reweighting on loss function for all baselines. The full details are included in Table VIII and VII in the Appendix.
@@ V-E Ablation Study
Our approach has two key pieces: a multi-objective optimization (MOO) paradigm and a margin-based class re-balancing strategy. Table VI reports an ablation with asynchronous inputs across three datasets. Both components consistently improve accuracy, and the best results arise when they are combined (bold). MOO is especially impactful when class imbalance is milder (FBG and WearGait), delivering gains of 12.58% and 12.57%, respectively, over the non-MOO counterpart. While each component alone outperforms the plain baseline, only their combination achieves the top performance across datasets (with the second-best underlined), highlighting the complementary nature of MOO and margin re-balancing.
Dataset MOO Margin Re-balancing Accuracy (%) FOG Dataset ✓ ✓ ✓ ✓ FBG Dataset ✓ ✓ ✓ ✓ WearGait Dataset ✓ ✓ ✓ ✓
@@ V-F Hyper-parameters Analysis
Additive Margin Coefficient .

### v2

@@ V-B Parkinson’s Disease Multimodal Datasets
We conducted evaluations on three publicly available multimodal PD gait datasets. All datasets contain gait data from two to three complementary modalities and accompanied by clinically annotated ratings. These characteristics make them suitable for the evaluation of our proposed framework.
FOG Dataset [52]: It contains video and IMU recordings from 35 PD subjects performing turning-in-place at 30 Hz (video) and 128 Hz (IMU) with clinical FoG annotations. Coordinates of 3D poses are obtained using MMPose [56] (MotionBERT trained on Human3.6M). FBG Dataset [53]: It comprises synchronized full-body kinematics and kinetics from 26 PD subjects walking overground in ON/OFF medication states, including 3D motion capture at 150 Hz and force-plate data, with MDS-UPDRS III and H&Y labels. WearGait Dataset [54]: It provides synchronized IMU, sensorized insole (16-sensor plantar pressures and embedded inertial signals), and walkway data, all at 100Hz, from 98 people with PD and 83 age-matched controls with clinical metadata.
@@ V-C Overall Evaluation
We evaluate the proposed TRIP framework under both asynchronous and synchronous conditions, each with and without class re-balancing, to comprehensively assess robustness and generalization. Single-modality oracles (e.g., Skeleton Only) serve as upper bounds, while quantifies how multimodal fusion compares to its strongest unimodal counterpart (negative indicates improvement).
Asynchronous Inputs. Under naturally imbalanced, asynchronous inputs—typical in clinical deployment—TRIP consistently achieves the highest accuracies (right blocks in Table III). For FOG, TRIP attains 56.68%, surpassing the best baseline (Late Fusion, 40.2%) and yielding a large multimodal gain (=41.97). Skeleton accuracy rises from 47.73% to 63.61% (+15.88 pp), and IMU from 33.47% to 49.76% (+16.29 pp). For FBG, it reaches 59.48%, outperforming DEEPAV (52.59%) by +6.89 pp, with Skeleton improved from 41.8% to 51.24% (+9.44 pp) and vGRF from 60.13% to 67.71% (+7.58 pp), resulting in =17.59. For WearGait, which includes three modalities, TRIP achieves 71.39%, outperforming all fusion methods while maintaining strong per-modality accuracies (Walkway 71.08%, Insole 63.03%, IMU 80.07%). These results confirm that TRIP can effectively learn from time-misaligned data without synchronized sampling, while other methods deteriorate.
Synchronous Inputs. When all modalities are synchronized (Table IV), TRIP retains its lead (right blocks). On FOG, it achieves 62.35%, outperforming all fusion baselines and exceeding the best unimodal oracle (57.49%) by +4.86 pp (=9.46). On WearGait, TRIP reaches 84.18%, the best among all models, with a substantial multimodal gain (=28.17). It surpasses all fusion competitors (e.g., Early Fusion 81.88%, Cross Attention 79.3%) and all single-modality oracles (Walkway 64.01%, Insole 59.59%, IMU 76.6%). We omit FBG synchronous results since inconsistent temporal alignment across sensor trials prevents fair synchronized evaluation, further underscoring the limitations of sync-dependent approaches.
Effect of Class Rebalancing. To confirm that improvements are not caused by label imbalance, we repeat experiments with inverse-frequency weighting on the loss function for all baseline methods. Under asynchronous + rebalanced training (left blocks in Table III), TRIP maintains its superiority in mean accuracy: 56.68% on FOG (=3.41), 59.48% on FBG (=2.72), and 71.39% on WearGait (=5.47). Under synchronous + rebalanced training (left blocks in Table IV), TRIP again performs best: 62.35% on FOG (=5.77) and 84.18% on WearGait (=25.46). Even when class frequencies are equalized, TRIP outperforms or matches the best rebalanced baselines (e.g., DEEPAV*, Early Fusion*), while preserving high per-modality accuracies.
Summary. Across all configurations—synchronous or asynchronous and balanced or unbalanced—TRIP consistently delivers the strongest fusion accuracy and the lowest . These findings confirm that: (1) TRIP robustly integrates complementary modalities without temporal alignment; (2) it enhances weaker modalities while maintaining strong ones via gradient-balanced optimization; and (3) it remains stable under class rebalancing, demonstrating intrinsic robustness to both modality and label imbalance.
@@ V-D Modality Collapse Mitigation
Dataset MOO Margin Rebalancing Accuracy (%) FOG Dataset ✓ ✓ ✓ ✓ FBG Dataset ✓ ✓ ✓ ✓ WearGait Dataset ✓ ✓ ✓ ✓
@@ V-E Ablation Study
Our approach has two key pieces: a multi-objective optimization (MOO) paradigm and a margin-based class rebalancing strategy. Table V reports an ablation with asynchronous inputs across three datasets. Both components consistently improve accuracy, and the best results arise when they are combined (bold). MOO is especially impactful when class imbalance is milder (FBG and WearGait), delivering gains of 12.58% and 12.57%, respectively, over the non-MOO counterpart. While each component alone outperforms the plain baseline, only their combination achieves the top performance across datasets (with the second-best underlined), highlighting the complementary nature of MOO and margin-based rebalancing.
@@ V-F Hyper-Parameters Analysis
Additive Margin Coefficient